A loaded boat weighs 981kg.The force exerted upon the boat in the direct of motion by the oars,is equivalent to a constant force of 15N and the resistance in Newton to the motion is equal to twice the speed of the boat . If th boat starts from rest, determine;

(i)the limiting speed

(ii)speed at the end of 10 sec.


1
Expert's answer
2021-09-28T11:41:58-0400

i)

"ma=15-2v"

"m\frac{dv}{dt}=15-2v"

"981\cdot\frac{dv}{dt}=15-2v"


"dt=\frac{981}{15-2v}dv"


"t=\int \frac{981}{15-2v}dv=-\frac{981}{2}ln(15-2v)+c"

"v=0" at "t=0"

"c=\frac{981}{2}ln15"


"t=-\frac{981}{2}ln(15-2v)+\frac{981}{2}ln15"


"e^{2t/981}=\frac{15}{15-2v}"


"v(t)=7.5-\frac{7.5}{e^{2t/981}}"


ii)

"v(10)=7.5-\frac{7.5}{e^{2\cdot 10/981}}=0.15" m/s